# Rings of non-commutative functions and their fields of fractions

### v1

@@ 3 Topologies on rings of analytic NC functions
Classically, the ring of holomorphic functions on a domain has a natural Fréchet topology that arises from a compact exhaustion of the domain. Every two such exhaustions give rise to the same topology. Moreover, the resulting topological algebra is locally multiplicatively convex (LMC) [32] and nuclear. One can view such algebras as inverse limits of function algebras. The quantized version of a function algebra is an operator algebra and a theory of quantized topological vector spaces was explored by Effros and Webster in [15, 53]. Later in Section 6 we will apply this theory of Effros and Webster to define a class of topological “local operator algebras” which admit well-defined evaluations of the universal skew field, $\mathscr{M}_{d}$, of $\mathscr{O}_{d}$.
@@
and where $\|f_{\ell}\|_{\mathbb{H}^{2}_{d}}$ is the norm of the $\ell$-th homogeneous component in $\mathbb{H}^{2}_{d}=\mathbb{H}^{2}_{d}(1)$. For any $r>0$, the Hardy algebra, $\mathbb{H}^{\infty}_{d}(r)$, defined above, can be identified with the left multiplier algebra of $\mathbb{H}^{2}_{d}(r)$, and $\mathbb{A}_{d}(r)$ can be equivalently defined as the operator-norm closure of the free algebra in $\mathbb{H}^{\infty}_{d}(r)$. The hardy algebra, $\mathbb{H}^{\infty}_{d}(r)$ is then the weak operator topology (WOT) closure of $\mathbb{A}_{d}(r)$ in $\mathscr{B}(\mathbb{H}^{2}_{d}(r))$. Also note that for any $0<r<s$, the re-scaling map, $F(Z)\mapsto F(\tfrac{r}{s}Z)$, defines an embedding of $\mathbb{H}^{\infty}_{d}(s)$ into $\mathbb{H}^{\infty}_{d}(r)$ and of $\mathbb{A}_{d}(s)$ in $\mathbb{A}_{d}(r)$, both with dense ranges (albeit, in the former case, in the weak* topology of $\mathscr{B}(\mathbb{H}^{2}_{d}(r))$). In particular, $F\in\mathbb{H}^{\infty}_{d}(r)$ if and only if $F(rL)\in\mathbb{H}^{\infty}_{d}=\mathbb{H}^{\infty}_{d}(1)$.
For $r<s$, there are natural restriction maps $\mathbb{A}_{d}(s)\hookrightarrow\mathbb{A}_{d}(r)$ and $\mathbb{H}^{2}_{d}(s)\hookrightarrow\mathbb{H}^{2}_{d}(r)$ which we will denote by $\pi_{r,s}$. This gives two families of topological rings, $\{\mathbb{A}_{d}(r)|\ r\in(0,\infty)\}$, and $\{\mathbb{H}^{2}_{d}(r)|\ r\in(0,\infty)\}$ equipped with continuous mappings $\pi_{r,s}$, $0<r<s<\infty$, given by restriction to the NC row-ball of radius $r<s$. For convenience, given any $R\in(0,+\infty]$, we fix a strictly increasing sequence $(r_{n})_{n=1}^{\infty}$ so that $\lim_{n\uparrow\infty}r_{n}=R$. In the case where $R=+\infty$ we will simply choose $r_{n}=n\in\mathbb{N}$. We will also write $\pi_{m,n}:=\pi_{r_{m},r_{n}}:\mathbb{A}_{d}(r_{n})\hookrightarrow\mathbb{A}_{d}(r_{m})$, $m<n$, for the continuous restriction map. Following [5], the inverse limit of $\{\mathbb{A}_{d}(r_{n})\}_{n=1}^{\infty}$ is then a topological ring defined as follows. First, consider the Cartesian product space
@@
We have two presentations of $\mathscr{O}_{d}$ as inverse limits
@@
The maps are the restriction maps defined above. Similarly,
[…]
Here, $0<r_{m}<R$ is any strictly increasing sequence with limit $R$.
@@
The first presentation is clear since the topology on $\mathscr{O}_{d}$ is given by the supremum norms on balls of increasing radius. For the second one, we note that if a function $f$ extends analytically beyond the ball, then there is a bound on the $\mathbb{H}^{\infty}_{d}$ norm by $\mathbb{H}^{2}_{d}$ norm, see Lemma 2.2. The same proof works for the claim on $\mathscr{O}_{d}(R)$, $R>0$. ∎
@@
It is easily checked that $\mathscr{O}_{d}(R)$ is complete with respect to the topology of uniform convergence on any NC sub-ball, $r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, $0<r<R$. ∎
Each of the algebras, $\mathscr{O}_{d}(R)$, satisfy a version of the classical Montel theorem. Though several Montel-type theorems exist in non-commutative analysis (see, for example [3, Theorem 13.14] and [41, Theorem 5.2]), we provide a slightly different proof here for the sake of completeness. Recall from [50, Definition 34.2] that a topological vector space $\mathcal{E}$ is called a Montel space, if it is Hausdorff, barrelled, and every closed bounded subset is compact. In particular, by [50, Corollary 3 in Section 50], every nuclear Fréchet space is Montel. In a locally convex space, boundedness is equivalent to boundedness in each of the seminorms defining the topology. In particular, for a Fréchet space a bounded sequence has a convergent subsequence by metrizability. Thus, we obtain a template for Montel-type theorems.
@@
By [46, 7.3 Corollary 3], a Fréchet space is nuclear, if it is an inverse limit of Hilbert spaces with trace-class connecting maps. Therefore, it suffices to show that the restriction map $\Phi_{r}\colon\mathbb{H}^{2}_{d}\to\mathbb{H}^{2}_{d}$ is trace-class. We note that since $\mathbb{H}^{2}_{d}=\oplus_{k=0}^{\infty}\mathbb{C}_{k}\langle\mathfrak{z}\rangle$, where $\mathbb{C}_{k}\langle\mathfrak{z}\rangle$ is the space of homogeneous free polynomials of degree $k$. On each $\mathbb{C}_{k}\langle\mathfrak{z}\rangle$, the map $\Phi_{r}$ acts as multiplication by $r^{k}$. Hence, $\Phi_{r}$ is trace class, provided that $r<d$, since $\mathrm{dim}\,\mathbb{C}_{k}\langle\mathfrak{z}\rangle=d^{k}$. Choose a cofinal system of radii and hence a cofinal system of semi-norms associated to NC row-balls of those radii, such that the radii grow exponentially, with exponent >d>d. It follows that $\mathscr{O}_{d}$ is the inverse limit of a system of Hilbert spaces with trace-class maps, as desired. ∎
@@
By [15, Theorem 7.4], $\mathscr{O}_{d}$ has a unique quantization as a local operator space. Later in 6 we will show that since $\mathscr{O}_{d}$ is an inverse limit of operator algebras, it can be viewed as a local operator algebra.
One can see here that the fact that we have an infinite radius of convergence is important in the proof. In fact, the algebras $S(r)$ introduced by Taylor, are not nuclear unless all but at most one of the radii are infinite [29]. It seems that the same reasoning applies in the row ball case. Therefore, we prove a weaker claim.
@@
By [18, I.6.2 Theorem], it suffices to show that if we have a set bounded in the norm induced from $\mathbb{H}_{2}^{d}(r_{m+1})$, then it contains a Cauchy subsequence in the norm from $\mathbb{H}^{2}_{d}(r_{m})$. However, this follows immediately from the fact that the restriction map from $\mathbb{H}_{2}^{d}(r_{m+1})$ to $\mathbb{H}^{2}_{d}(r_{m})$ is compact. Indeed, setting $t:=\tfrac{r_{m}}{r_{m+1}}<1$, we can identify the restriction map with the linear re-scaling map, $\Phi_{t}\in\mathscr{B}(\mathbb{H}^{2}_{d})$ defined by $\Phi_{t}(f):=f(tL)1$. It is clear that this linear map is diagonal with respect to the standard orthonormal basis of monomials for $\mathbb{H}^{2}_{d}$, and $\Phi_{t}(\mathfrak{z}^{\omega})=t^{|\omega|}\mathfrak{z}^{\omega}$. Hence the diagonal entries of $\Phi_{t}$ converge to 0 as $|\omega|\rightarrow\infty$ so that $\Phi_{t}$, and hence the restriction map, are compact. Since a bounded subset in $\mathbb{H}_{2}^{d}(r_{m+1})$ is weakly precompact, its image in $\mathbb{H}_{2}^{d}(r_{m})$ is pre-compact in norm. ∎
Classically, the ring of entire functions is a Bézout domain. (And a commutative ring is a semifir if and only if it is a Bézout domain.) Moreover, every finitely generated ideal of the ring of entire functions (and even the ring of holomorphic functions on a domain in $\mathbb{C}$) is principal and closed. This result is called Helmer’s theorem according to [28, Theorem 13.6]. The converse also holds. Namely, if an ideal of the ring of entire functions is closed, then it is principal. The key tool in the proof of those results is the Weierstrass theorem that allows one to build entire functions with prescribed zeroes and multiplicities. In other words, the closed ideal of the ring of entire functions are described by analytic varieties (counting the multiplicity). This leads us to define varieties in the NC entire setting. Varieties of various types were studied both in the algebraic setting [21] and the setting of the NC unit row ball [24, 43]. Let $S\subseteq\mathscr{O}_{d}$. Define the corresponding variety similarly to [24],
@@
Let $g\notin\overline{J}$. Then, there exists a continuous functional $\varphi$ on $\mathscr{O}_{d}(R)$, such that $\varphi(g)\neq 0$, but $\varphi|_{\overline{J}}=0$. Since $\mathscr{O}_{d}(R)$ is an inverse limit, there exists $m\in\mathbb{N}$, such that $\varphi$ factors through $\mathbb{H}^{2}_{d}(r_{m})$. Therefore, there is an $h\in\mathbb{H}^{2}_{d}(r_{m})$, such that for all $f\in J$ and every free polynomial $p$, $\langle{h},{M^{R}_{p}f}\rangle_{\mathbb{H}^{2}_{d}(r_{m})}=0$. Therefore, $h$ is orthogonal to the range of $f(r_{m}L)$. However, by our assumption. $\langle{h},{g}\rangle_{\mathbb{H}^{2}_{d}(r_{m})}\neq 0$. Hence, $g\notin I(\mathscr{Z}(J))$. ∎
@@ 4 Algebraic properties of rings of analytic NC functions
In this section, we will prove two main results. First, we will prove that each $\mathscr{O}_{d}(R)$ is a semifir and we will also show that every finitely–generated right ideal in $\mathscr{O}_{d}(R)$ is closed. From the latter fact, we derive a version of the Bergman Nullstellensatz for each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$. Recall from [10] that a ring, $\mathscr{R}$, is called a semifir or semi-free ideal ring if every finitely generated right (equivalently, left) ideal is free as a right (respectively, left) module. Alternative conditions for $\mathscr{R}$ to be a semifir are provided in [10, Theorem 2.3.1]. In particular, we will employ the condition [10, Theorem 2.3.1(a)]. In order to state this condition, recall that for two $n$-tuples of elements $x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}\in\mathscr{R}$, if $\sum_{j=1}^{n}x_{j}y_{j}=0$, we say that this equation is an $n$-term relation for $\mathscr{R}$. Such a relation is said to be trivial, if for every $1\leq i\leq n$, either $x_{i}=0$ or $y_{i}=0$. Write $x=\left(\begin{smallmatrix}x_{1}\\ \vdots\\ x_{n}\end{smallmatrix}\right)\in\mathscr{R}^{n}$ and $y=\left(\begin{smallmatrix}y_{1}\\ \vdots\\ y_{n}\end{smallmatrix}\right)\in\mathscr{R}^{n}$. We say that an invertible matrix $S\in\mathscr{R}^{n\times n}$ trivializes the relation $x^{\mathrm{t}}y=0$, if the new relation $x^{{}^{\prime};\mathrm{t}}y^{\prime}:=(x^{\mathrm{t}}S)(S^{-1}y)=0$ is trivial. We can now state the condition [10, Theorem 2.3.1(a)]; a ring, $\mathscr{R}$, is a semifir if and only if every finite relation, $\sum_{j=1}^{n}x_{j}y_{j}=0$, can be trivialized by an invertible matrix in $\mathscr{R}^{n\times n}$. Since our goal is to prove that $\mathscr{O}_{d}(R)$ is a semifir, for any $0<R\leq+\infty$, we will fix an arbitrary $n\in\mathbb{N}$ and elements $f_{1},\ldots,f_{n},g_{1},\ldots,g_{n}\in\mathscr{O}_{d}(R)$, such that $\sum_{j=1}^{n}f_{j}g_{j}=0$. We will write
$F=\begin{pmatrix}f_{1}&\ldots&f_{n}\end{pmatrix}\in\mathscr{O}_{d}(R)^{1\times n}\text{ and }G=\begin{pmatrix}g_{1}\\ \vdots\\ g_{n}\end{pmatrix}\in\mathscr{O}_{d}(R)^{n}.$
Then, our relation is $FG=0$. Our basic strategy will be to first show that given any such $n$-term relation for $\mathscr{O}_{d}(R)$, it can be trivialized in $\mathbb{A}_{d}(r)$ for any $0<r<R$. We then construct an inverse limit of invertible $n\times n$ matrices over $\mathbb{A}_{d}(r)$, $0<r<R$ which trivialize our relation in $r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, i.e. in $\mathbb{A}_{d}(r)$, to obtain an invertible matrix which trivializes our original relation for the inverse limit ring, $\mathscr{O}_{d}(R)$.
@@
Let $\Phi\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{1\times n}$ . Then, there exists $S\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$, such that $\Phi S=(\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{k},0,\cdots,0)$. The row, $\tilde{\Phi}=(\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{k})$, is injective if and only if there exists a decomposition $\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}=\mathcal{H}\oplus\mathrm{Ker}\,\Phi$, where $\mathcal{H}$ is a closed right shift-invariant subspace. Moreover, given any other $T\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$, so that $\Phi T$ has a decomposition with its first $k$-coordinates an injective row and the last nkn-k components are 0, we have that $T$ satisfies the same properties as $S$ and
$T=S\begin{pmatrix}A&0\\ B&C\end{pmatrix}$.
@@
If there is such a decomposition, then we can choose $S=(V_{1},V_{2})$, where $V_{1}$ is the inner with range $\mathcal{H}$ and $V_{2}$ is the inner with range $\mathrm{Ker}\,\Phi$. Conversely, if there is an $S$, such that $\Phi S$ has the form above, then set $W=\mathrm{Span}\{e_{n-\ell},\ldots,e_{n}\}\otimes\mathbb{H}^{2}_{d}$ and note that $S^{-1}\mathrm{Ker}\,\Phi=W$. Therefore, if $\mathcal{H}=SW^{\perp}$, then $\mathcal{H}+\mathrm{Ker}\,\Phi=\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}$. If $T$ is another such invertible matrix then $T^{-1}SW=W$. Hence, $T^{-1}S$ is lower-triangular and by a symmetric argument, so is $S^{-1}T$. ∎
@@
Clearly, this relation holds in $\mathbb{A}_{d}(r)$, i.e. for $S^{-1}(\tfrac{r}{s}L)G(rL)$. Lastly,

### v2

@@ 3 Topologies on rings of analytic NC functions
Classically, the ring of holomorphic functions on a domain in $\mathbb{C}^{d}$ has a natural Fréchet topology that arises from a compact exhaustion of the domain. Every two such exhaustions give rise to the same topology. Moreover, the resulting topological algebra is locally multiplicatively convex (LMC) [33] and nuclear. One can view such algebras as inverse limits of function algebras. The quantized version of a function algebra is an operator algebra and a theory of quantized topological vector spaces was explored by Effros and Webster in [16, 55]. Later in Section 6 we will apply this theory of Effros and Webster to define a class of topological “local operator algebras” which admit well-defined evaluations of the universal skew field, $\mathscr{M}_{d}$, of $\mathscr{O}_{d}$.
@@
and where $\|f_{\ell}\|_{\mathbb{H}^{2}_{d}}$ is the norm of the $\ell$-th homogeneous component in $\mathbb{H}^{2}_{d}=\mathbb{H}^{2}_{d}(1)$. For any $r>0$, the Hardy algebra, $\mathbb{H}^{\infty}_{d}(r)$, defined above, can be identified with the left multiplier algebra of $\mathbb{H}^{2}_{d}(r)$, and $\mathbb{A}_{d}(r)$ can be equivalently defined as the operator–norm closure of the free algebra in $\mathbb{H}^{\infty}_{d}(r)$. The Hardy algebra, $\mathbb{H}^{\infty}_{d}(r)$ is then the weak operator topology (WOT) closure of $\mathbb{A}_{d}(r)$ in $\mathscr{B}(\mathbb{H}^{2}_{d}(r))$. Also note that for any $0<r<s$, the re-scaling map, $F(Z)\mapsto F(\tfrac{r}{s}Z)$, defines an embedding of $\mathbb{H}^{\infty}_{d}(s)$ into $\mathbb{H}^{\infty}_{d}(r)$ and of $\mathbb{A}_{d}(s)$ in $\mathbb{A}_{d}(r)$, both with dense ranges. (Albeit, in the former case, in the weak* topology of $\mathscr{B}(\mathbb{H}^{2}_{d}(r))$.) In particular, $F\in\mathbb{H}^{\infty}_{d}(r)$ if and only if $F(rL)\in\mathbb{H}^{\infty}_{d}=\mathbb{H}^{\infty}_{d}(1)$.
For $r<s$, there are natural restriction maps $\mathbb{A}_{d}(s)\hookrightarrow\mathbb{A}_{d}(r)$ and $\mathbb{H}^{2}_{d}(s)\hookrightarrow\mathbb{H}^{2}_{d}(r)$ which we will denote by $\pi_{r,s}$. This gives two families of topological rings, $\{\mathbb{A}_{d}(r)|\ r\in(0,\infty)\}$, and $\{\mathbb{H}^{2}_{d}(r)|\ r\in(0,\infty)\}$ equipped with continuous mappings $\pi_{r,s}$, $0<r<s<\infty$, given by restriction to the NC row-ball of radius $r<s$. For convenience, given any $R\in(0,+\infty]$, we fix a strictly increasing sequence $(r_{n})_{n=1}^{\infty}$ of positive radii so that $\lim_{n\uparrow\infty}r_{n}=R$. In the case where $R=+\infty$ we will simply choose $r_{n}=n\in\mathbb{N}$. We will also write $\pi_{m,n}:=\pi_{r_{m},r_{n}}:\mathbb{A}_{d}(r_{n})\hookrightarrow\mathbb{A}_{d}(r_{m})$, $m<n$, for the continuous restriction map. Following [5], the inverse limit of $\{\mathbb{A}_{d}(r_{n})\}_{n=1}^{\infty}$ is then a topological ring defined as follows. First, consider the Cartesian product space
@@
For any $R\in(0,+\infty]$, we have two presentations of $\mathscr{O}_{d}(R)$ as inverse limits
@@
The maps are the restriction maps defined above, and $0<r_{m}<R$ is any strictly increasing sequence with limit $R$.
@@
Consider $R=+\infty$. The first presentation is clear since the topology on $\mathscr{O}_{d}$ is given by the supremum norms on balls of increasing radius. For the second one, we note that if a function $f$ extends analytically beyond the ball, then there is a bound on the $\mathbb{H}^{\infty}_{d}$ norm by $\mathbb{H}^{2}_{d}$ norm, see Lemma 2.2. The same proof works for the claim on $\mathscr{O}_{d}(R)$, $R>0$. ∎
@@
It is easily checked that $\mathscr{O}_{d}(R)$ is complete with respect to the topology of uniform convergence on every NC sub-ball, $r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, $0<r<R$. ∎
Each of the algebras, $\mathscr{O}_{d}(R)$, satisfy a version of the classical Montel theorem. Though several Montel-type theorems exist in non-commutative analysis (see, for example [3, Theorem 13.14] and [43, Theorem 5.2]), we provide a slightly different proof here for the sake of completeness. Recall from [52, Definition 34.2] that a topological vector space $\mathcal{E}$ is called a Montel space, if it is Hausdorff, barrelled, and every closed bounded subset is compact. By [52, Corollary 3 in Section 50], every nuclear Fréchet space is Montel. In a locally convex space, boundedness is equivalent to boundedness in each of the seminorms defining the topology. In particular, for a Fréchet space a bounded sequence has a convergent subsequence by metrizability. Thus, we obtain a template for Montel-type theorems.
@@
[…]
@@
By [16, Theorem 7.4], $\mathscr{O}_{d}$ has a unique quantization as a local operator space. Later in 6 we will show that since $\mathscr{O}_{d}$ is an inverse limit of operator algebras, it can be viewed as a local operator algebra.
One can see here that the fact that we have an infinite radius of convergence for any element of $\mathscr{O}_{d}$ is important in the proof of Theorem 3.3. In fact, the algebras $S(r)$ introduced by Taylor (see Remark 2.1), are not nuclear unless all but at most one of the radii are infinite [30]. It seems that the same reasoning applies in our case of $\mathscr{O}_{d}(r)$, $0<r<+\infty$, and so we prove a weaker result for these rings.
@@
By [19, I.6.2 Theorem], it suffices to show that if we have a set bounded in the norm induced from $\mathbb{H}_{2}^{d}(r_{m+1})$, then it contains a Cauchy subsequence in the norm from $\mathbb{H}^{2}_{d}(r_{m})$. However, this follows immediately from the fact that the restriction map from $\mathbb{H}_{2}^{d}(r_{m+1})$ to $\mathbb{H}^{2}_{d}(r_{m})$ is compact. Indeed, setting $t:=\tfrac{r_{m}}{r_{m+1}}<1$, we can identify the restriction map with the linear re-scaling map, $\Phi_{t}\in\mathscr{B}(\mathbb{H}^{2}_{d})$ defined by $\Phi_{t}(f):=f(tL)1$. It is clear that this linear map is diagonal with respect to the standard orthonormal basis of monomials for $\mathbb{H}^{2}_{d}$, and $\Phi_{t}(\mathfrak{z}^{\omega})=t^{|\omega|}\mathfrak{z}^{\omega}$. Hence the diagonal entries of $\Phi_{t}$ converge to 0 as $|\omega|\rightarrow\infty$ so that $\Phi_{t}$, and hence the restriction map, are compact. Since a bounded subset in $\mathbb{H}_{2}^{d}(r_{m+1})$ is weakly precompact, its image in $\mathbb{H}_{2}^{d}(r_{m})$ (under the restriction map, $\pi_{m,m+1}$) is pre-compact in norm. ∎
The following lemma will be useful in the next section. It describes the density of the image of invertible elements under the restriction map.
Lemma 3.6.
Let $0<r<s$. For every $n\in\mathbb{N}$, the image of $\operatorname{GL}_{n}(\mathcal{O}_{d})$ in $\operatorname{GL}_{n}(\mathbb{A}_{d}(r))$ is dense. In particular, the image of $\operatorname{GL}_{n}(\mathbb{A}_{d}(s))$ in $\operatorname{GL}_{n}(\mathbb{A}_{d}(r))$ is dense.
Proof.
Let us fix $F\in\operatorname{GL}_{n}(\mathbb{A}_{d}(r))$. Let us consider the norm continuous path $\gamma\colon[0,1]\to\operatorname{GL}_{n}(\mathbb{A}_{d}(r))$ given by $\gamma(t)(Z)=F((1-t)Z)$, for every $Z\in r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. Clearly, $\gamma$ connects $F$ to the constant function $F(0)$. Using the fact that $\operatorname{GL}_{n}(\mathbb{C})$ is path connected, we connect $F(0)$ to the identity matrix. By a known characterization of the path-connected component to the identity in the group of invertible elements of a unital Banach algebra, there exist $G_{1},\ldots,G_{m}\in M_{n}(\mathbb{A}_{d}(r))$, such that $F=e^{G_{1}}\cdot e^{G_{2}}\cdots e^{G_{m}}$. (See, for example, [15, Theorem 2.14].) Using Cesàro sums (as discussed in the proof of Lemma 2.3), we find sequences of matrices of polynomials $Q_{k}^{(j)}\rightarrow G_{j}$ in $\mathbb{A}_{d}(r)$. It is straightforward to check that the exponential map on a Banach algebra is continuous. (This can be done directly, or by applying upper semi-continuity of the spectrum and the Riesz–Dunford holomorphic functional calculus.) Hence, the sequence of NC entires $H_{k}=e^{Q_{k}^{(1)}}\cdot e^{Q_{k}^{(2)}}\cdots e^{Q_{k}^{(m)}}$ converges to $F$ in $\operatorname{GL}_{n}(\mathbb{A}_{d}(r))$. ∎
Classically, the ring of entire functions is a Bézout domain. (And a commutative ring is a semifir if and only if it is a Bézout domain.) Moreover, every finitely generated ideal of the ring of entire functions (and even the ring of holomorphic functions on a domain in $\mathbb{C}$) is principal and closed. This result is called Helmer’s theorem according to [29, Theorem 13.6] and the converse also holds. Namely, if an ideal of the ring of entire functions is closed, then it is principal. The key tool in the proof of those results is the Weierstrass product theorem that allows one to build entire functions with prescribed zeroes and multiplicities. In other words, closed ideals in the ring of entire functions are described by analytic varieties (counting the multiplicity). Varieties of various types have been considered in the non-commutative setting [22, 25, 45]. In what follows we consider an “NC directional variety”, in the sense of G. M. Bergman, see [24]. Given $S\subseteq\mathscr{O}_{d}$, define the corresponding variety as in [25],
@@
Let $g\notin\overline{J}$. Then, there exists a continuous functional $\varphi$ on $\mathscr{O}_{d}(R)$, such that $\varphi(g)\neq 0$, but $\varphi|_{\overline{J}}=0$. Since $\mathscr{O}_{d}(R)$ is an inverse limit, there exists $m\in\mathbb{N}$, such that $\varphi$ factors through $\mathbb{H}^{2}_{d}(r_{m})$. Therefore, there is an $h\in\mathbb{H}^{2}_{d}(r_{m})$, such that for all $f\in J$ and every free polynomial $p$, $\langle{h},{M^{R}_{p}f}\rangle_{\mathbb{H}^{2}_{d}(r_{m})}=0$ and hence $h$ is orthogonal to the range of $f(r_{m}L)$. However, by our assumption, $\langle{h},{g}\rangle_{\mathbb{H}^{2}_{d}(r_{m})}\neq 0$ so that $g\notin I(\mathscr{Z}(J))$. ∎
@@ 4 Algebraic properties of rings of analytic NC functions
In this section, we will prove two main results. First, we will prove that each $\mathscr{O}_{d}(R)$ is a semifir, and secondly we will show that every finitely–generated right ideal in $\mathscr{O}_{d}(R)$ is closed. From the latter fact, we derive a version of the Bergman Nullstellensatz for each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$. Recall from [10] that a ring, $\mathscr{R}$, is called a semifir or semi-free ideal ring if every finitely generated right (equivalently, left) ideal is free as a right (respectively, left) module. Alternative conditions for $\mathscr{R}$ to be a semifir are provided in [10, Theorem 2.3.1]. In particular, we will employ the condition [10, Theorem 2.3.1(a)]. In order to state this condition, recall that for two $n$-tuples of elements $x_{1},\cdots,x_{n},y_{1},\cdots,y_{n}\in\mathscr{R}$, if $\sum_{j=1}^{n}x_{j}y_{j}=0$, this equation is said to be an $n$-term relation for $\mathscr{R}$. Such a relation is said to be trivial, if for every $1\leq i\leq n$, either $x_{i}=0$ or $y_{i}=0$. Write $x=\left(\begin{smallmatrix}x_{1}\\ \vdots\\ x_{n}\end{smallmatrix}\right)\in\mathscr{R}^{n}$ and $y=\left(\begin{smallmatrix}y_{1}\\ \vdots\\ y_{n}\end{smallmatrix}\right)\in\mathscr{R}^{n}$. We say that an invertible matrix $S\in\mathscr{R}^{n\times n}$ trivializes the relation $x^{\mathrm{t}}y=0$, $x^{\mathrm{t}}:=(x_{1},\cdots,x_{n})\in\mathscr{R}^{1\times n}$, if the new relation $x^{{}^{\prime};\mathrm{t}}y^{\prime}:=(x^{\mathrm{t}}S)(S^{-1}y)=0$ is trivial. We can now state the condition [10, Theorem 2.3.1(a)]: A ring, $\mathscr{R}$, is a semifir if and only if every finite relation, $\sum_{j=1}^{n}x_{j}y_{j}=0$, $n\in\mathbb{N}$, can be trivialized by an invertible matrix in $\mathscr{R}^{n\times n}$. Since our goal is to prove that $\mathscr{O}_{d}(R)$ is a semifir, for any $0<R\leq+\infty$, we will fix an arbitrary $n\in\mathbb{N}$ and elements $f_{1},\cdots,f_{n},g_{1},\cdots,g_{n}\in\mathscr{O}_{d}(R)$, such that $\sum_{j=1}^{n}f_{j}g_{j}=0$. We will write
$F=\begin{pmatrix}f_{1}&\cdots&f_{n}\end{pmatrix}\in\mathscr{O}_{d}(R)^{1\times n}\qquad\text{and}\qquad G=\begin{pmatrix}g_{1}\\ \vdots\\ g_{n}\end{pmatrix}\in\mathscr{O}_{d}(R)^{n},$
so that our relation is $F\cdot G=0$. Our basic strategy will be to first show that given any such $n$-term relation for $\mathscr{O}_{d}(R)$, it can be trivialized in $\mathbb{A}_{d}(r)$ for any $0<r<R$. We then construct an inverse limit of invertible $n\times n$ matrices over $\mathbb{A}_{d}(r)$, $0<r<R$, which trivialize our relation in $r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, i.e. in $\mathbb{A}_{d}(r)$, to obtain an invertible matrix over $\mathscr{O}_{d}(R)$ that trivializes our original relation for the inverse limit ring, $\mathscr{O}_{d}(R)$.
@@
[…]
$T=S\begin{pmatrix}A&0\\ B&C\end{pmatrix};\qquad\qquad A\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{(n-\ell)\times(n-\ell)},\ B\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{\ell\times(n-\ell)},\ C\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{\ell\times\ell}.$
@@
If there is such a decomposition, then we can choose $S=(V_{1},V_{2})$, where $V_{1}$ is the inner with range $\mathcal{H}$ and $V_{2}$ is the inner with range $\mathrm{Ker}\,\Phi$. Conversely, if there is an $S$, such that $\Phi S$ has the form above, then set $\mathcal{W}=\bigvee\{e_{n-\ell+1},\cdots,e_{n}\}\otimes\mathbb{H}^{2}_{d}$ and note that $S^{-1}\mathrm{Ker}\,\Phi=\mathcal{W}$. Therefore, if $\mathcal{H}=S\mathcal{W}^{\perp}$, then $\mathcal{H}\dotplus\mathrm{Ker}\,\Phi=\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}$. If $T$ is another such invertible matrix then $T^{-1}S\mathcal{W}=\mathcal{W}$. Hence, $T^{-1}S$ is block lower-triangular and by a symmetric argument, so is $S^{-1}T$. ∎
@@
Clearly, this relation holds in $\mathbb{A}_{d}(r)$, i.e. for $S(\tfrac{r}{s}L)^{-1}G(rL)$. Lastly,